OpenStudy (anonymous):

Gustav is using front-end estimation. In estimating seventy-two point nine three minus thirty-two point zero seven, Gustav follows the steps below. Where did he go wrong? Step 1: seventy-two minus thirty-two equals forty Step 2: zero point nine three minus zero point zero seven is approximately equal to one Step 3:forty minus one equals thirty-nine Step 4: seventy-two point nine three minus thirty-two point zero seven is approximately equal to thirty-nine Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

72.93 - 32.07 72 - 32 = 40 0.93 - 0.07 = .86 (close to 1) 40 - 1 = 39 72.93 - 32.07 = 40.86 (it approximately equal to 40, not 39) step 4 is incorrect